### v1

# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

Abstract
Knockoffs are a powerful tool for controlled variable selection with false discovery rate (FDR) control. However, while they are frequently used in high-dimensional regressions, they lack power in low-dimensional and sparse signal settings. One of the main reasons is that knockoffs require a minimum number of selections, depending on the nominal FDR level. In this paper, we leverage e-values to allow the nominal level to be switched after looking at the data and applying the knockoff procedure. In this way, we can increase the nominal level in cases where the original knockoff procedure does not make any selections to potentially make discoveries. Also, in cases where the original knockoff procedure makes discoveries, we can often decrease the nominal level to increase the precision. These improvements come without any costs, meaning the results of our post-hoc knockoff procedure are always more informative than the results of the original knockoff procedure. Furthermore, we apply our technique to recently proposed derandomized knockoff procedures.
1 Introduction
In exploratory analyses, the identification of relevant variables is a critical step toward understanding complex data-generating mechanisms and uncovering meaningful associations. [^1] We refer to the task as variable selection, noting that equivalent terminology includes feature selection, biomarker discovery, or covariate identification, depending on the application domain. Over the past years, a wide range of methods that leverage modern machine learning techniques have been proposed to quantify the importance of candidate variables. While these approaches provide powerful tools for ranking variables according to their apparent influence, they typically lack rigorous guarantees for controlled variable selection. In particular, translating variable importance scores into a selected subset of variables while controlling false discoveries remains a major methodological challenge. To bridge this gap, recent methods have aimed to develop inference procedures that combine statistical rigor with modeling flexibility, with the knockoff framework emerging as a particularly effective approach [1, 4].
1.1 Controlled variable selection and knockoffs
In this work, we consider the problem of variable selection under rigorous false discovery control. Suppose we observe a response $Y$ and variables $\mathbf{X}=(X_{1},\dots,X_{p})$. For each variable $X_{i}$, we consider the conditional independence null hypothesis
$H_{0}^{i}:X_{i}\perp\!\!\!\perp Y\mid\mathbf{X}_{-i},$ (2)
where $\mathbf{X}_{-i}$ denotes all variables except $X_{i}$. Variables for which $H_{0}^{i}$ holds are called null, and the remaining are non-null (or relevant). We denote these sets by
$I_{0}=\{i:H_{0}^{i}\text{ is true}\},\qquad I_{1}=[p]\setminus I_{0}.$ (3)
A variable selection procedure outputs a set of discoveries/rejections $R\subseteq[p]$, and its quality is measured by the false discovery rate
$\mathrm{FDR}=\mathbb{E}\!\left[\frac{|R\cap I_{0}|}{|R|\vee 1}\right],$ (4)
where $a\vee b=\max(a,b)$ avoids division by zero.
The task of identifying the relevant variables, that is, those in $I_{1}$, naturally leads to the question of how to design a variable selection procedure that can recover as many non-nulls as possible while rigorously controlling false discoveries. The knockoff framework of Barber and Candès [1], refined and generalized by Candès et al. [4], addresses this goal by enabling controlled variable selection: it yields a data-driven rejection set of putative discoveries while provably controlling the false discovery rate (FDR) at a user-specified level $\alpha\in(0,1)$, meaning $\textnormal{FDR}\leq\alpha$. Methods based on knockoffs construct synthetic variables that act as negative controls, enabling data-adaptive selection with provable FDR guarantees, while making minimal assumptions on the conditional distribution of $Y\mid\mathbf{X}$. A full description of the knockoff construction and filtering procedure is deferred to Section 2.1.
1.2 Motivation of our work
While the knockoff framework was primarily developed for high-dimensional problems, it is equally valuable to have methods that provide controlled variable selection in low- and moderate-dimensional settings, particularly when the underlying signal is expected to be sparse. In many scientific applications, the number of truly relevant variables is small, even if a large number of variables are measured. When identifying effect modifiers (i.e., predictive variables) in clinical trials, the total number of available clinical variables is typically modest—often on the order of 10–50 patient characteristics—reflecting the scale of routinely collected baseline data. Within this set, it is generally assumed that only a small subset truly drives heterogeneity in treatment effect, leading to a sparse predictive structure in most methodological developments and simulation studies [16, 24, 17, 30]. Interestingly, in such settings, the vanilla knockoff procedure exhibits certain limitations that motivate further methodological development. In our work we will focus on two main limitations of the original knockoffs:
The minimum number of discoveries required to achieve statistical significance is determined by the pre-specified FDR level $\alpha$, and it is $1/\alpha,$ which can be overly conservative when only a few signals are present. As a result, the procedure may fail to flag any discoveries even when a small but genuine signal exists.
If the difficulty of the problem (e.g., signal strength or sparsity) is not well understood beforehand — which is not rare in exploratory analyses, especially when new data are being studied — choosing the FDR level $\alpha$ a priori can be challenging. Once the analysis is conducted, there is no valid post-hoc adjustment of $\alpha$ to adapt to the observed evidence, which may lead to either low power or weak error guarantee.
These limitations underscore the need for methods that maintain the rigorous error control of the knockoff framework while providing greater flexibility and improved sensitivity in sparse or low-signal settings. In the remainder of this paper, we develop and assess such approaches, demonstrating how post-hoc adjustment of the nominal FDR level can overcome these challenges.
1.3 Contribution
Our paper addresses the aforementioned two key limitations of standard FDR-controlling knockoff methods by leveraging e-values to adjust the nominal FDR level after observing the data.
Switching $\alpha$ post-hoc with knockoffs:
E-values [19, 20] are a novel tool for hypothesis testing that has proven particularly useful for sequential anytime-valid inference [7, 27, 34, 9], testing of composite nulls in irregular models [33, 15], multiple testing [32, 11, 31, 36], and, most recently, for hypothesis testing at data-dependent nominal levels [8, 14, 6, 36]. In this paper, we exploit the two latter aspects of e-values to construct a knockoff procedure that allows to switch the nominal FDR level post-hoc (see Section 2.2 for a detailed description of FDR control with data-dependent $\alpha$). This improves the original knockoff method [1, 4] in two ways.
  1. 1.
    If the original knockoff method does not make any discoveries, we can increase the nominal level to potentially obtain rejections.
  2. 2.
    If the original knockoff method makes discoveries, we can often decrease the nominal level while keeping the same discovery set, thus increasing the precision of the results.
Figure 1 illustrates the benefits of the proposed post-hoc adjustment. In Figure 1(a), we observe that the post-hoc procedure dramatically increases power: across all settings for the number of relevant variables, our method successfully identifies nearly all relevant variables (i.e., power $\approx 1$). In contrast, the original knockoff procedure exhibits very low power when the number of relevant variables is small. This is primarily because the minimum number of discoveries is determined by the nominal FDR level $\alpha$, which in our case corresponds to $1/\alpha=5$. To illustrate this, when the actual number of relevant variables is 3, the original procedure produces no discoveries in 1542 out of 2000 iterations and exactly 5 discoveries in 229 iterations, which are the two most frequent outcomes. By contrast, the post-hoc method most frequently selects 3 variables (1002 iterations) and 4 variables (467 iterations), corresponding to the two most common outcomes, closely matching the true number of relevant variables.
The gains in power come with only a small sacrifice in the (averaged) nominal $\alpha$ value, as illustrated in Figure 1(b) for scenarios with a low number of relevant variables (3 and 4). Moreover, our post-hoc adjustment can reduce the effective nominal level $\alpha$ when the problem is easier than anticipated, while maintaining the same power as the original procedure. It should be noted that our improvement is a “free-lunch”, meaning our approach never makes less discoveries and it only reports a larger significance level $\alpha$ in cases where the original method did not make any rejections. Therefore, the results reported by our method are always more informative.
(a) Power vs number of relevant variables
(b) FDR and $\alpha$ vs number of relevant variables
Figure 1: This figure illustrates the main contribution of our work, namely that the proposed post-hoc adjustment of the $\alpha$ values substantially increases statistical power compared to the original approach. For small values of $p_{\text{relevant}}$, this improvement comes at a minor cost of slightly inflated $\alpha$ values, whereas for larger $p_{\text{relevant}}$, the post-hoc method can even reduce $\alpha$ while maintaining comparable power. The nominal level of $\alpha$ for the original KO method is set to $0.20.$ A full description of the data-generating process, parameter choices, and simulation protocol is provided in Section 5.
Post-hoc $\alpha$ with derandomized knockoffs:
Since the knockoff procedure is based on the random generation of knockoff variables, the results are heavily exposed to external randomness. In order reduce its dependence on the random data generator, Ren et al. [22], Ren and Barber [21] introduced derandomized knockoff procedures by sampling multiple knockoff copies for each variable. These derandomized knockoff procedures require the pre-specification of an additional parameter, on which the performance and guarantee of the procedure depend considerably. Ren et al. [22], Ren and Barber [21] proposed heuristics, simulations and data-splitting methods to obtain a reasonable parameter value. We show that the parameter can be chosen arbitrarily based on the entire data, thus increasing the efficiency and utility of the derandomized methods enormously. Furthermore, we show how the derandomized procedures can be further uniformly improved.
1.4 Outline
In the following, we start to briefly summarize the knockoff procedure by Barber and Candès [1], Candès et al. [4] and the e-Closure Principle to derive FDR procedures with post-hoc error control by Xu et al. [36]. Afterwards, in Section 3, we introduce our post-hoc knockoff procedure and derive its error control. In Section 4, we consider the derandomized procedures and our proposed improvements. Finally, in Section 5, we provide extensive simulation results that illustrate the superiority of our methods. [^2] The code implementing the methods described in this paper is publicly available at https://github.com/sechidis/posthoc_knockoffs.
2 Preliminaries
2.1 The BC-knockoff procedure
The knockoff procedure introduced by Barber and Candès [1] and extended by Candès et al. [4] is a powerful framework for variable selection while controlling the false discovery rate (FDR). Since we consider multiple variations of their initial approach in this paper, we call it BC-knockoff procedure in the following. The strategy is composed of three steps:
  1. 1.
    Generate Knockoff Variables: For each input variable $X_{i}$, an artificial variable $\tilde{X}_{i}$ called a knockoff is generated to act as a negative control for assessing the importance of the original variable. To ensure the knockoff variables can serve this purpose, they are constructed with the following two critical properties:
    Property (a) - Exchangeability.
[…]
    Property (b) - Conditional Independence.
    Given the original variables $\mathbf{X}$, the knockoffs $\tilde{\mathbf{X}}$ are conditionally independent of the response variable $Y$. That is:
    $\tilde{\mathbf{X}}\perp\!\!\!\perp Y\,|\,\mathbf{X}.$ (5)
    This property is straightforwardly guaranteed as long as neither the response $Y$ in the knockoff generation process. Its purpose is to maintain neutrality during model fitting and selection steps.
    Various algorithms have been proposed to construct knockoffs that obey the exchangability property [26, 23, 13, 2, 29, 10].
  2. 2.
    Estimate Knockoff Statistics: For each variable $X_{i}$, we calculate knockoff statistics $W_{i}$ to measure how much more important the original variable is compared to its knockoff $\tilde{X}_{i}$. A positive value of $W_{i}$ indicates that the original variable is more important, where the magnitude captures the strength of the evidence. The derivation of knockoff statistics follows a two-step procedure:
    Step 2a - Calculate Variable Importance Scores.
    Using the dataset that includes both original variables $\mathbf{X}$ and their knockoffs $\tilde{\mathbf{X}}$, along with the outcome variable $Y$, variable importance scores are computed for both original and knockoff variables. Specifically, we calculate a statistic of statistics (Z,Z~)($\tilde{Z}$), representing the importance measures for each pair of original and knockoff variables:
    $(Z,\tilde{Z})=(Z_{1},\dots,Z_{p},\tilde{Z}_{1},\dots,\tilde{Z}_{p})=q\left\{(\mathbf{X},\tilde{\mathbf{X}}),Y\right\},$
    where $Z_{i}$ and $\tilde{Z}_{i}$ capture importance for the original variable $X_{i}$ and its knockoff $\tilde{X}_{i}$, respectively. The function $q$ represents an appropriate model or statistical procedure that assigns importance scores. These statistics are constructed so that swapping any variable with its knockoff results in the corresponding components of the statistic being swapped as well. Most variable importance score estimation procedures, such as those based on regression coefficients and permutation importance measures, inherently satisfy this symmetry property [4]. This exchangeability is a key requirement for maintaining the validity of knockoffs.
    Step 2b - Derive Knockoff Statistics.
    Using the variable importance scores from Step 2a, the knockoff statistic $W_{i}$ is defined as:
    $W_{i}=f(Z_{i},\tilde{Z}_{i}),$
    where $f$ is an antisymmetric function satisfying f(u,v)=f(v,u)f(u,v)=-f(v,u). This property ensures that swapping $X_{i}$ with its knockoff $\tilde{X}_{i}$ flips the sign of $W_{i}$. Common choices for $f$ include:
    $f(u,v)=u-v,\quad\text{or}\quad f(u,v)=|u|-|v|.$ (6)
    A large, positive $W_{i}$ provides evidence against the null hypothesis that $X_{i}$ is an unimportant variable, allowing it to be included in the rejection set in Step 3.
    The last years, many powerful knockoff statistics have been introduced in the literature, see for example Spector and Fithian [28]. Furthermore, there are works that derive knockoff statistics targeting effect modification in randomized trials: for example, Sechidis et al. [25] develop predictive knockoff statistics that quantify the extent to which baseline variables modify treatment effects in RCTs, enabling FDR-controlled selection of predictive variables (i.e., effect modifiers) rather than purely prognostic ones.
  3. 3.
    Threshold Knockoff Statistics: Let $\epsilon_{i}$ be the sign of $W_{i}$ and $\mathcal{W}=\sigma(|W_{1}|,\ldots,|W_{p}|)$. Define the rejection set $R^{\textnormal{BC}}$ for a predefined level $\alpha^{\textnormal{kn}}\in(0,1]$ as:
    $R^{\textnormal{BC}}\coloneqq\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}\},\quad\text{where }T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}\coloneqq\inf\left\{t>0:\frac{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}}\geq 1/\alpha^{\textnormal{kn}}\right\}.$ (7)
    If there does not exist a $t>0$ satisfying the above condition, then $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$. The key theoretical result that forms the foundation of this methodology is stated in this Lemma:
    Fact 1 (Lemma 3.3 in [4]).
    Conditional on $\mathcal{W}$, the signs corresponding to true hypotheses $(\epsilon_{i})_{i\in I_{0}}$ are i.i.d. coin flips.
With the latter result, one can show that the BC-knockoff procedure controls the FDR at level $\alpha=\alpha^{\textnormal{kn}}$ [1, 4]. Similar to [21], we distinguish between the FDR level $\alpha$ and the parameter $\alpha^{\textnormal{kn}}$ used for the knockoff stopping time in (7), since they can deviate later in the paper. Also, in this paper we propose to choose the significance level $\alpha$ post-hoc, however, the parameter $\alpha^{\textnormal{kn}}$ needs to be specified in advance and can be interpreted as an initial or targeted significance level.
2.2 Post-hoc choice of $\alpha$
The Closure Principle of Marcus et al. [18] has long been known as a fundamental approach to multiple testing with familywise error rate (FWER) control. Recently, Xu et al. [36] introduced an e-Closure Principle for false discovery rate (FDR) control. Given a local e-value $E_{S}$ for each intersection hypothesis $H_{0}^{S}=\bigcap_{i\in S}H_{0}^{i}$, $S\subseteq[p]$, meaning $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, we consider the closure set
$\mathcal{R}_{\alpha}(\boldsymbol{E})\coloneqq\left\{R\subseteq[p]:E_{S}\geq\frac{\textnormal{FDP}_{S}(R)}{\alpha}\text{ for all }S\subseteq[p]\right\},$ (8)
where $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$ and $\textnormal{FDP}_{S}(R)=\frac{|S\cap R|}{|R|\lor 1}$. Xu et al. [36] showed that any rejection set $R\in\mathcal{R}_{\alpha}(\boldsymbol{E})$ controls the FDR and any procedure that controls the FDR at level $\alpha$ can be constructed in this way.
Furthermore, Xu et al. [36] showed that these closure sets even permit a more sophisticated error rate control that goes beyond the classical notion of FDR control — $\alpha$ can be chosen post-hoc! Given a family of e-values $\boldsymbol{E}=(E_{S})_{S\subseteq[p]}$, they proved that
$\mathbb{E}\left[\sup_{\alpha\in(0,1]}\max_{R\in\mathcal{R}_{\alpha}(\boldsymbol{E})}\frac{\textnormal{FDP}_{I_{0}}(R)}{\alpha}\right]\leq 1.$ (9)
This particularly implies that
$\mathbb{E}\left[\frac{\textnormal{FDP}_{I_{0}}(\tilde{R})}{\tilde{\alpha}}\right]\leq 1\text{ for any data-dependent }\tilde{\alpha}\in(0,1],\tilde{R}\in\mathcal{R}_{\tilde{\alpha}}(\boldsymbol{E}).$ (10)
Therefore, after calculating the e-values $\boldsymbol{E}$, we can construct the closure sets $\mathcal{R}_{\alpha}(\boldsymbol{E})$ for multiple $\alpha$’s and then decide completely data-dependent which $\tilde{\alpha}\in(0,1]$ and $\tilde{R}\in\mathcal{R}_{\tilde{\alpha}}(\boldsymbol{E})$ to report while guaranteeing (10).
Note that the post-hoc error rate control in (10) is slightly different from classical FDR control (where the FDR is bounded by $\alpha$), since $\tilde{\alpha}$ is inside the expectation. If $\tilde{\alpha}$ was fixed in advance, then the two notions would be equivalent due to the linearity of the expected value. However, since $\tilde{\alpha}$ is allowed to be a random variable depending on the data, the classical notion needs to be adapted by drawing $\tilde{\alpha}$ inside the expectation. This notion of post-hoc validity was also considered in [8, 14] for single hypotheses.
[…]
Example 1.
Suppose all null hypotheses are true ($I_{0}=[p]$). In that case $\textnormal{FDP}_{I_{0}}(R)$ becomes $1$ if $R\neq\emptyset$ and 0 otherwise. Suppose we have a multiple testing procedure $R_{\alpha}$ for each $\alpha\in(0,1]$ that has exact error rate control, meaning $\mathbb{E}[\textnormal{FDP}_{I_{0}}(R_{\alpha})]=\alpha$ (for example, this is satisfied by the Benjamini-Hochberg procedure [3] with independent and uniformly distributed null p-values). Now set $\tilde{\alpha}=\inf\{\alpha\in(0,1]:R_{\alpha}\neq\emptyset]$. Then, we have
$\mathbbm{E}\left[\frac{\textnormal{FDP}_{I_{0}}\left(R_{\tilde{\alpha}}\right)}{\tilde{\alpha}}\right]=\int_{0}^{1}1/x\ dx=\infty.$
Therefore, cherry-picking the significance level with fixed-level procedures can inflate the expected value in (10) indefinitely.
In this paper, we will develop a knockoff procedure that gains power by allowing the significance level to be chosen post-hoc, while ensuring that (9) (and thus (10)) is fulfilled.
3 Switching $\alpha$ post-hoc with knockoffs
Considering the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ in (7), it is apparent that at least $1/\alpha^{\textnormal{kn}}=1/\alpha$ variables with a strong signal are needed to make rejections with BC-knockoff, as $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$ otherwise. For this reason, BC-knockoff often performs poorly in low dimensional settings (see Figure 1). In this section, we show how we can exploit the e-Closure Principle (Section 2.2) in such cases to increase the significance level $\alpha$ (but not the parameter $\alpha^{\textnormal{kn}}$) post-hoc and potentially make some rejections while guaranteeing FDR control in the sense of (10). Furthermore, we also show that in cases where BC-knockoff makes rejections, we can often decrease the significance level, thus increasing the precision.
3.1 Increase $\alpha$ to gain power
First, let us define the stopping time
$T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\coloneqq\inf\left\{t>0:\frac{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}}\geq 1/\alpha^{\textnormal{kn}}\quad\text{or}\quad\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0\right\},$ (11)
where the parameter $\alpha^{\textnormal{kn}}\in(0,1]$ can be interpreted as the prespecified initial significance level. The difference of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ of Barber and Candès [1] (see equation (7)) is the second condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$. Be aware that this condition only comes into play if the first condition is not satisfied for any $t$, meaning $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ if $R^{\textnormal{BC}}\neq\emptyset$. However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we have $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}<\infty$. Similar to BC-knockoff, our procedure, which we call post-hoc knockoff (ph-knockoff), rejects all hypotheses $H_{0}^{i}$ with $W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$. Since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$, ph-knockoff is uniformly more powerful than BC-knockoff by Barber and Candès [1]. However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we can only report the rejections at a larger significance level $\alpha>\alpha^{\textnormal{kn}}$, as we show in the next section (see Theorem 3.1).
Remark 1.
Ren and Barber [21] introduced a modification of the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ that is similar to $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$, adding the condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}<1/\alpha^{\textnormal{kn}}$ instead of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$ (see Remark 5 in [21]). They used this adjusted stopping time to improve their derandomized knockoff procedure (see also Section 4) without considering a post-hoc choice of $\alpha$. While our approach of switching $\alpha$ post-hoc would also work with their stopping time, we think our stopping time leads to smaller post-hoc levels.
3.2 Decrease $\alpha$ to improve precision
Up to this point, we have only considered to choose a level $\alpha>\alpha^{\textnormal{kn}}$ to potentially obtain more rejections. However, it turns out that in cases where BC-knockoff makes rejections ($R^{\textnormal{BC}}\neq\emptyset$), it actually controls the FDR at a smaller level $\alpha$ than the prespecified $\alpha^{\textnormal{kn}}$. Since rejections at a smaller significance level indicate more evidence against the null hypothesis, we can decrease $\alpha$ compared to the initial significance level $\alpha^{\textnormal{kn}}$ post-hoc to improve the precision of the results.
In Algorithm 10, we summarize our general procedure, where we denote by $\alpha^{\textnormal{ph}}$ our post-hoc significance level and by $R^{\textnormal{ph}}$ the corresponding rejection set. If BC-knockoff does not make any rejections, we increase the significance level ($\alpha^{\textnormal{ph}}>\alpha^{\textnormal{kn}}$), and thus potentially obtain rejections. If BC-knockoff makes rejections, we decrease the significance level ($\alpha^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$) to increase the evidence. In this way, the claims made by ph-knockoff are always stronger than the claims made by BC-knockoff.
Algorithm 1 ph-knockoff
Input: Initial significance level $\alpha^{\textnormal{kn}}$ and knockoff statistics $W_{1},\ldots,W_{p}$.
Output: Rejection set $R^{\textnormal{ph}}$ and post-hoc level $\alpha^{\textnormal{ph}}$.
1:Calculate $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ from (11)
[…]
3:if $R^{\textnormal{ph}}\neq\emptyset$ then
4:  $\alpha^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}$
5:else
6:  $\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$
7:end if
8:return $R^{\textnormal{ph}}$, $\alpha^{\textnormal{ph}}$
To prove the guarantee of our procedure, we define the local e-values
$E_{S}^{\textnormal{ph}}=\frac{\sum_{i\in S}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}},S\subseteq[p].$ (12)
The validity of these local e-values (i.e. that $\mathbb{E}_{H_{0}^{S}}[E_{S}^{\textnormal{ph}}]\leq 1$) follows since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ is a stopping time with respect to the filtration generated by a masked version of $(W_{i})_{i\in[p]}$, meaning the event $\{T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq t\}$ is determined by the absolute values $|W_{i}|$ for all $i$ and $\sum_{i:|W_{i}|\geq s}\mathbbm{1}\{W_{i}>0\}$ for all $s\leq t$ (see Remark 4 in [21]).
Theorem 3.1.
Let $\boldsymbol{E}^{\textnormal{ph}}=(E_{S}^{\textnormal{ph}})_{S\subseteq[p]}$ be the family of local e-values defined in (12) and set
$\alpha^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}\quad\text{with the convention }\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}\text{ if }\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0.$ (13)
Then $R^{\textnormal{ph}}\in\mathcal{R}_{\alpha^{\textnormal{ph}}}(\boldsymbol{E}^{\textnormal{ph}})$, where $R^{\textnormal{ph}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}$. Therefore, it holds that
$\mathbb{E}\left[\frac{\textnormal{FDP}_{I_{0}}(R^{\textnormal{ph}})}{\alpha^{\textnormal{ph}}}\right]\leq 1.$ (14)
Furthermore, $R^{\textnormal{ph}}\supseteq R^{\textnormal{BC}}$, meaning ph-knockoff always makes as least as many rejections as BC-knockoff and $\alpha^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$ on the event $\{R^{\textnormal{BC}}\neq\emptyset\}$, meaning if BC-knockoff rejects any hypothesis then ph-knockoff does not report a larger significance level.
Proof.
We need to show that $E_{S}^{\textnormal{ph}}\geq\textnormal{FDP}_{S}(R^{\textnormal{ph}})/\alpha^{\textnormal{ph}}$ for all $S\subseteq[p]$. First, note that in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0$, we have $R^{\textnormal{ph}}=\emptyset$ and therefore the condition is trivially fulfilled. Hence, assume $R^{\textnormal{ph}}\neq\emptyset$. Then,
$E_{S}^{\textnormal{ph}}=\frac{|S\cap R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})|R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})}{\alpha^{\textnormal{ph}}}.$ (15)
Remark 2.
Note that $\alpha^{\textnormal{ph}}$ equals 11 in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=1$. While rejections at a level of 11 seems like a trivial claim in the fixed-$\alpha$ regime, it is not for random $\alpha$ [14]. However, if such large values for $\alpha^{\textnormal{ph}}$ should be avoided, one could just set $\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$ and $R^{\textnormal{ph}}=\emptyset$ in those cases. This obviously maintains (14).
4 Post-hoc $\alpha$ with derandomized knockoffs
Ren and Barber [21] introduced a derandomized knockoff procedure, which we call RB-knockoff in the following, by generating multiple knockoff copies and merging the results using an average of (compound) e-values. In the following, we briefly recap their method and demonstrate how it can be improved by using a post-hoc significance level. It should be noted that we present the RB-knockoff procedure with our stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to be consistent with the previous part, even though Ren and Barber [21] proposed a slightly different one (see Remark 1). However, our approach generally works for any stopping time.
Let $W_{i}^{(1)},\ldots,W_{i}^{(k)}$, $i\in[p]$, be $k$ knockoff statistics for the $i$-th variable. It is only required that the vector $W^{(j)}=(W_{1}^{(j)},\ldots,W_{p}^{(j)})$ satisfies the property in Fact 1 for each $j\in[k]$, but the dependence structure between the different $W^{(j)}$ can be arbitrary. Now let $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$ be the stopping time (11) applied on $W^{(j)}$ and define for each $i\in[p]$,
$E_{i}^{\text{avg}}=\frac{1}{k}\sum_{j=1}^{k}E_{i}^{(j)},\quad\text{where }E_{i}^{(j)}=p\frac{\mathbbm{1}\{W_{i}^{(j)}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}\}}{1+\sum_{l\in[p]}\mathbbm{1}\{W_{l}^{(j)}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}\}}.$ (16)
Ren and Barber [21] showed that $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ are compound e-values, meaning $\mathbb{E}\left[\sum_{i\in I_{0}}E_{i}^{\text{avg}}\right]\leq p$. The RB-knockoff procedure is defined by e-BH [32] applied on $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ at prespecified level $\alpha^{\textnormal{ebh}}\in(0,1]$ and controls the FDR at level $\alpha^{\textnormal{ebh}}$.
In case of $k=1$ and $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$, RB-knockoff coincides with BC-knockoff [21]. However, Ren and Barber [21] showed that $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$ often leads to zero power in case of $k>1$ and therefore recommend to choose $\alpha^{\textnormal{ebh}}>\alpha^{\textnormal{kn}}$. The choice of the parameter $\alpha^{\textnormal{ebh}}$ is crucial. If $\alpha^{\textnormal{ebh}}$ is too small, one might not be able to make any rejections even though the signal is strong. If $\alpha^{\textnormal{ebh}}$ is too large, the results become imprecise. Ren and Barber [21] derive some concrete choices like $\alpha^{\textnormal{ebh}}=2\alpha^{\textnormal{kn}}$ based on simulations and heuristics, but they do not propose an optimal parameter. We note that it is possible to choose $\alpha^{\textnormal{ebh}}$ completely data-dependent while maintaining (10). In this way, we can just choose the $\alpha^{\textnormal{ebh}}$ that leads to the most desirable rejection set.
Theorem 4.1.
Let $E_{(1)}^{\text{avg}}\geq\ldots\geq E_{(p)}^{\text{avg}}$ be the ordered compound e-values defined in (16) and $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\{i\in[p]:E_{i}^{\text{avg}}\geq p/(\alpha^{\textnormal{ebh}}i^{\textnormal{ebh}})\}$, where $i^{\textnormal{ebh}}=\max\{i\in[p]:E_{(i)}^{\text{avg}}\geq p/(\alpha^{\textnormal{ebh}}i)\}$ with the convention $\max(\emptyset)=0$. Then,
$\mathbb{E}\left[\sup_{\alpha^{\textnormal{ebh}}\in(0,1]}\frac{\textnormal{FDP}_{I_{0}}(R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}})}{\alpha^{\textnormal{ebh}}}\right]\leq 1.$ (17)
Hence, the parameter $\alpha^{\textnormal{ebh}}$ can be chosen based on the data while maintaining FDR control in the sense of (10).
The proof of the Theorem 4.1 follows immediately from the equivalence of e-BH with compound e-values $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ and the e-Closure Principle with the following local e-values (see Theorem 23 in [36])
$E_{S}=\frac{1}{p}\sum_{i\in S}E_{i}^{\text{avg}}.$ (18)
There are multiple reasonable choices to pick $\alpha^{\textnormal{ebh}}$ based on the data that could lead to different rejection sets (in the following we refer to our data-dependent choice as $\alpha^{\textnormal{dph}}$). In order to compare our derandomized ph-knockoff procedure to RB-knockoff in the following, we start with some initial level $\alpha^{\textnormal{ebh}}\in(0,1)$ that should be chosen as in the RB-knockoff procedure. If RB-knockoff does not make any rejections with that $\alpha^{\textnormal{ebh}}$, we choose $\alpha^{\textnormal{dph}}>\alpha^{\textnormal{ebh}}$ as small as possible such that $R_{\alpha^{\textnormal{dph}}}^{\textnormal{RB}}$ is not empty. If $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then we choose $\alpha^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$ as small as possible while $R_{\alpha^{\textnormal{dph}}}^{\textnormal{RB}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}$. In this way, our derandomized ph-knockoff procedure always rejects as least as many hypotheses as RB-knockoff. Furthermore, if RB-knockoff makes rejections, then $\alpha^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$. Our general procedure is summarized in Algorithm 2.
Algorithm 2 Derandomized ph-knockoff
Input: Knockoff level $\alpha^{\textnormal{kn}}\in(0,1]$, initial level $\alpha_{\textnormal{ebh}}\in[0,1]$ and knockoff statistics $W_{1}^{(j)},\ldots,W_{p}^{(j)}$ for each $j\in[k]$.
Output: Rejection set $R^{\textnormal{dph}}$ and post-hoc level $\alpha^{\textnormal{dph}}$.
1:Calculate $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ as in (16)
2:Set $i^{\textnormal{ebh}}=\max\{i\in[p]:E_{(i)}^{\text{avg}}\geq p/(\alpha^{\textnormal{ebh}}i)\}$
3:Define the RB-knockoff rejection set $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\{i\in[p]:E_{i}^{\text{avg}}\geq p/(\alpha^{\textnormal{ebh}}i^{\textnormal{ebh}})\}$
4:if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$ then
5:  $\alpha^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})$
6:  $R^{\textnormal{dph}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}$
7:else
8:  $i^{*}=\operatorname*{argmax}_{i\in[p]}iE_{(i)}^{\text{avg}}$ (take largest index if maximum is not unique)
9:  if $i^{*}E_{(i^{*})}^{\text{avg}}<p$ then
10:   $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{kn}}$
11:   $R^{\textnormal{dph}}=\emptyset$
12:  else
13:   $\alpha^{\textnormal{dph}}=p/(i^{*}E_{(i^{*})}^{\text{avg}})$
14:   $R^{\textnormal{dph}}=\{i\in[p]:E_{i}^{\text{avg}}\geq p/(i^{*}\alpha^{\textnormal{dph}})\}$
15:  end if
16:end if
17:return $R^{\textnormal{dph}}$, $\alpha^{\textnormal{dph}}$
Remark 3.
To get rid of the (arbitrary) parameter $\alpha^{\textnormal{ebh}}$ in Algorithm 2, one could set it to default values. For example, if $\alpha^{\textnormal{ebh}}=0$, then the derandomized ph-knockoff procedure chooses the smallest $\alpha^{\textnormal{dph}}\in(0,1]$ that leads to at least one rejection (and sets $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{kn}}$ if no such $\alpha^{\textnormal{dph}}$ exists). Furthermore, if $\alpha^{\textnormal{ebh}}=1$, then Algorithm 2 chooses the smallest $\alpha^{\textnormal{dph}}\in(0,1]$ that leads to the same rejection set as RB-knockoff applied at level $\alpha^{\textnormal{ebh}}=1$.
Remark 4.
In case of $k=1$ the derandomized ph-knockoff method (Algorithm 2) and the ph-knockoff procedure (Algorithm 1) are equivalent regardless of the chosen $\alpha_{\textnormal{ebh}}\in[0,1]$. To see this, note that for $k=1$ we have that
$E_{i}^{\text{avg}}=p\frac{\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{1+\sum_{l\in[p]}\mathbbm{1}\{W_{l}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}.$
Hence, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then $R^{\textnormal{dph}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=R^{\textnormal{ph}}$ and $\alpha^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}=\alpha^{\textnormal{ph}}$. Furthermore, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\emptyset$, it holds that $i^{*}=\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}$, which implies $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{ph}}$ and $R^{\textnormal{dph}}=R^{\textnormal{ph}}$ in the same manner as before. Hence, the post-hoc framework unifies the non-derandomized and the derandomized knockoff method.
4.1 A uniform improvement of the derandomized (ph-)knockoff procedure
In this paragraph, we introduce a uniform improvement of the derandomized knockoff procedure for fixed $\alpha^{\textnormal{ebh}}$ by Ren and Barber [21] and therefore also of the derandomized ph-knockoff procedure. However, this uniform improvement is based on the e-Closure Principle and we have not yet found a short-cut. This makes this uniform improvement computationally inefficient, which is why we still recommend Algorithm 2 for practical purposes.
To derive the uniform improvement, consider the local e-values $(E_{S})_{S\subseteq[p]}$ that recover the RB-knockoff procedure, defined in (18). One can write $E_{S}$, $S\subseteq[p]$, as
$E_{S}=\frac{1}{k}\left(\sum_{j=1}^{k}\frac{\sum_{i\in S}\mathbbm{1}\{W_{i}^{(j)}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}^{(j)}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}\right).$ (19)
Now, due to the proof for FDR control of BC-knockoff by Barber and Candès [1], it follows that one can also define local e-values by changing $[p]$ in the denominators to $S$:
$\tilde{E}_{S}=\frac{1}{k}\left(\sum_{j=1}^{k}\frac{\sum_{i\in S}\mathbbm{1}\{W_{i}^{(j)}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}{1+\sum_{i\in S}\mathbbm{1}\{W_{i}^{(j)}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph},(j)}\}}\right).$ (20)
Obviously, $\tilde{E}_{S}\geq E_{S}$ for all $S\subseteq[p]$ and thus the e-Closure Principle with the local e-values $(\tilde{E}_{S})_{S\subseteq[p]}$, called closed knockoff in the following, never rejects less hypotheses than the RB-knockoff procedure. This was already noted by Xu et al. [36] for $k=1$ (knockoff without derandomization). However, while they showed that in case of $k=1$ closed knockoff is just equivalent to BC-knockoff, it becomes a real uniform improvement for $k>1$ as we illustrate in the following example.
Example 2.
[…]
4.2 Derandomized knockoffs for PFER control
Ren et al. [22] introduced a derandomized version of a knockoff procedure [12] for PFER control. We call this procedure RWC-knockoff in the following. The PFER is defined as the expected number of false discoveries $\textnormal{PFER}=\mathbb{E}[|R\cap I_{0}|]=\mathbb{E}[\textnormal{NFD}_{I_{0}}(R)]$, where $\textnormal{NFD}_{I_{0}}(R)=|R\cap I_{0}|$ for some rejection set $R$.
Let $\nu>0$ be an integer and $T_{\nu}^{\text{PFER}}=\inf\{t>0:\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=\nu-1\}$ be the stopping time for the PFER knockoff procedure introduced by Janson and Su [12]. Again, we denote by $T_{\nu}^{\text{PFER},(j)}$, $j\in[k]$, the stopping time in the $j$-th knockoff run. Then RWC-knockoff is given by the rejection set
$R_{\eta}^{\textnormal{RWC}}=\left\{i\in[p]:\left(\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right)\geq\eta\right\},$ (21)
where $\eta\in(0,1]$ is some prespecified parameter that determines after how many “rejections” in the different knockoff runs a hypothesis is finally rejected. Ren et al. [22] showed that RWC-knockoff controls the PFER at level $\nu/\eta$. Again, there is no optimal choice for $\eta$ and they propose sample splitting to obtain a data-driven parameter. We show that $\eta$ can be chosen entirely based on the data while retaining a sensible guarantee.
Theorem 4.2.
For some prespecified integer $\nu>0$ and the rejection set $R_{\eta}^{\textnormal{RWC}}$ in (21) it holds that
$\mathbb{E}\left[\sup_{\eta\in(0,1]}\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})\eta\right]\leq\nu.$ (22)
Hence, we can choose $\eta$ based on the data while maintaining $\mathbb{E}[\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})\eta]\leq\nu$.
Proof.
For the proof we implicitly use the e-Closure Principle for PFER control by Xu et al. [36]. Define the local e-values $E_{S}$, $S\subseteq[p]$, as in (18) but with the stopping times $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$. To see that $\mathbb{E}_{H_{0}^{S}}[E_{S}]\leq 1$, see for example Remark 4 in [21]. For any $\eta\in(0,1]$ it holds that
$\frac{\eta}{\nu}\textnormal{NFD}_{I_{0}}(R_{\eta}^{\textnormal{RWC}})$ $=\frac{\eta}{\nu}\sum_{i\in I_{0}}\mathbbm{1}\left\{\left(\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}\right)\geq\eta\right\}$ (23)
$\leq\frac{1}{\nu}\sum_{i\in I_{0}}\frac{1}{k}\sum_{j=1}^{k}\mathbbm{1}\{W_{i}^{(j)}\geq T_{\nu}^{\text{PFER},(j)}\}=E_{I_{0}},$ (24)
where the inequality follows from $\mathbbm{1}\{x\geq 1\}\leq x$ for all $x\geq 0$. Since $\mathbb{E}[E_{I_{0}}]\leq 1$, the claim follows. ∎
In this case we recommend to just calculate the rejection set $R_{\eta}^{\textnormal{RWC}}$ for each $\eta\in\{r/k:r\in[k]\}$ and then choose the combination of $\eta$ and $R_{\eta}^{\textnormal{RWC}}$ that fits best.
Remark 5.
In the same way as in Section 4.1 we can also uniformly improve RWC-knockoff by using $(\tilde{E}_{S})_{S\subseteq[p]}$ in (20) with $T_{\nu}^{\text{PFER},(j)}$ instead of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph},(j)}$ and then apply the e-Closure Principle for PFER control by Xu et al. [36].
5 Simulated experiments
5.1 Generating data
Our simulation framework builds on the setup of Ren and Barber [21]. As in the original work, we simulate data from both a Gaussian linear model and a logistic regression model. To reflect the two simulation regimes explored in Section 5.3—low-dimensional and high-dimensional—we describe each separately below.
Across all experiments, the variable vector $\mathbf{X}\in\mathbb{R}^{p}$ is generated as
$\mathbf{X}\sim\mathcal{N}(0,\Sigma),\qquad\Sigma_{jk}=0.5^{|j-k|}.$ (25)
The response conditional on $\mathbf{X}$ is drawn from either
$\textnormal{Gaussian model: }Y\mid\mathbf{X}\sim\mathcal{N}(\mathbf{X}^{\top}\beta,1),\qquad\textnormal{Logistic model: }Y\mid\mathbf{X}\sim\textnormal{Bernoulli}\!\left(\frac{e^{\mathbf{X}^{\top}\beta}}{1+e^{\mathbf{X}^{\top}\beta}}\right).$ (26)
Construction of $\beta$.
To define the regression coefficients, we first sample a signal vector $\bar{\beta}\in\mathbb{R}^{p_{\text{relevant}}}$ with i.i.d. entries from $\mathcal{N}(A,1)$, where $A$ controls signal amplitude. The full vector $\beta\in\mathbb{R}^{p}$ is obtained by embedding these non-zero entries into $p$ coordinates at approximately equally spaced locations, alternating their signs and scaling them by $\sqrt{n}$:
$\beta=\big(\underbrace{0,\ldots,0}_{z},\tfrac{\bar{\beta}_{1}}{\sqrt{n}},\underbrace{0,\ldots,0}_{z},-\tfrac{\bar{\beta}_{2}}{\sqrt{n}},\ldots\big).$ (27)
The spacing parameter $z$ is chosen as a function of $p$ and $p_{\text{relevant}}$, ensuring that signals are distributed evenly throughout the variable vector. As in Ren and Barber [21], this alternating-sign, evenly spaced construction reduces correlation between active variables while maintaining signal heterogeneity.
Low-dimensional regime (Section 5.3.1).
To investigate settings where classical knockoffs experience loss in power, we first consider the low-dimensional case, where
$p=50,\qquad n=250,\qquad p_{\text{relevant}}\in\{1,\ldots,10\}.$ (28)
For each choice of $p_{\text{relevant}}$, the spacing $z$ is adapted to distribute non-null variables evenly. We vary the signal amplitude on the grid
$A\in\{2,\dots,10\}\text{ (Gaussian)},\qquad A\in\{8,\dots,16\}\text{ (Logistic)}.$ (29)
Results for this regime are presented in Figures 2 and 3, where we compare post-hoc knockoffs to the original knockoff procedure of Candès et al. [4].
High-dimensional regime (Section 5.3.2).
To enable direct comparison with the derandomized knockoff procedure of Ren and Barber [21], we also simulate $n=1000$ samples in the same high-dimensional settings, namely
$p=800\text{ (Gaussian)},\qquad p=600\text{ (Logistic)}.$ (30)
The original study considered $p_{\text{relevant}}=80$ with $z=9$ for the Gaussian model and $p_{\text{relevant}}=60$ with $z=11$ for the logistic model. We adopt the same parameterization, and further study sparser cases by varying
$p_{\text{relevant}}\in\{20,40,80\}\text{ (Gaussian)},\qquad p_{\text{relevant}}\in\{20,40,60\}\text{ (Logistic)}.$ (31)
The signal amplitude ranges match those used in the low-dimensional study above. Results for this regime appear in Figures 4 and 5, where we benchmark our derandomized post-hoc procedure against the original derandomized knockoffs method.
5.2 Evaluation measures
For each model and each signal amplitude setting, the dataset $\mathcal{D}=(\mathbf{X},Y)$ is generated independently 100 times. For each draw of the dataset, we apply the knockoffs procedures that we are comparing. For each model, each signal amplitude, and each of the two methods, we measure performance in two different ways. First, we estimate the power and FDR as
$\textnormal{Power}=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{1}|}{|I_{1}|}\quad\textnormal{and}\quad\textnormal{FDR}=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1},$ (32)
where $R_{d}$ denotes the rejected set for the $d$-th dataset $\mathcal{D}_{d}$.
[…]
$\textnormal{Average ratio FDP/}\alpha=\frac{1}{100}\sum_{d=1}^{100}\frac{|R_{d}\cap I_{0}|}{|R_{d}|\vee 1}\frac{1}{\alpha_{d}},$ (33)
where $\alpha_{d}$ is the significance level $\alpha$ obtained in the $d$-th simulation run.
5.3 Results
5.3.1 Comparing with the original knockoffs in low-dimensional settings
In this section we illustrate the performance of our proposed post-hoc knockoff (presented in Algorithm 1) and compare to the original knockoff procedure from Candès et al. [4]. The details of the simulation setup are presented in Section 5.1. Through the simulation we set the $\alpha^{\textnormal{kn}}$ level to $0.20$ FDR, i.e., this is the nominal level of the traditional knockoffs, and the initial significance level of our post-hoc knockoffs.
In Figure 2, we evaluate the performance of the different methods as a function of $p_{\text{relevant}}$, with signal strength held fixed. When the number of truly relevant variables is small, traditional methods suffer from very low power. In contrast, post-hoc knockoffs offer the flexibility to substantially increase power, with paying some price in the probability of committing a Type-I error. For $p_{\text{relevant}}>3$, the increase in Type-I error is small relative to the power gain. Importantly, across nearly all settings, the empirical FDR remains close to the nominal target level $\alpha$, comparable to that of the original knockoff procedure. Furthermore, when power is comparable across methods (e.g., Gaussian setting for $p_{\text{relevant}}>6$), the post-hoc approach can achieve stricter error control, resulting in effective $\alpha$ values less than $0.20$.
Similar behavior is observed when $p_{\text{relevant}}$ is held constant and the signal amplitude is varied, as shown in Figure 3. Post-hoc knockoffs consistently achieve higher power than traditional methods, especially at moderate signal strengths, with the price of increasing $\alpha$ above $0.20$. For high signal amplitudes, our method achieves high power, and at the same time, it can provide more stringent control, i.e. $\alpha<0.20.$ Interestingly, the empirical FDR remains below the nominal $\alpha$ level in all scenarios.
Gaussian
Logistic
[…]
Gaussian
Logistic
[…]
5.3.2 Comparing with the derandomized knockoffs in high-dimensional settings
In this section, we illustrate the performance of our proposed derandomized post-hoc knockoff procedure (presented in Algorithm 2) and compare it with the original derandomized knockoff procedure from Ren and Barber [21]. The details of the simulation setup are presented in Section 5.1. For the nominal levels, we adopt the configurations from Ren and Barber [21]: we set the FDR level $\alpha=0.20$, which corresponds to the $\alpha_{\textnormal{ebh}}$ level of the derandomized knockoffs and serves as the initial significance level in our post-hoc derandomized knockoffs procedure, while in both methods we set $\alpha^{\textnormal{kn}}=\frac{\alpha_{\textnormal{ebh}}}{2}.$
Figure 4 presents the results under the Gaussian linear model, where the dimensionality is $p=800$. The last column, corresponding to $p_{\text{relevant}}=80$, reflects the setting considered in Ren and Barber [21]. In this regime, both methods exhibit comparable performance overall. However, our post-hoc procedure yields a modest power gain at lower signal amplitudes, albeit with a slight increase in the Type I error rate. This trade-off becomes more pronounced in sparser settings, where $p_{\text{relevant}}=20$ and $40$. Notably, in highly sparse scenarios, our method can also achieve a lower Type I error rate while maintaining power comparable to the original derandomized approach. For example, when $p_{\text{relevant}}=20$ (first column) and the signal amplitude exceeds $5$, our procedure attains similar power with improved Type I error control.
Figure 5 presents the results under the logistic regression model, where the dimensionality is $p=600$. The final column, corresponding to $p_{\text{relevant}}=60$, serves as the primary reference setting. In this regime, both methods demonstrate overall comparable performance. Nevertheless, our post-hoc procedure provides a moderate power improvement at lower signal amplitudes, accompanied by a slight increase in the Type I error rate. This trade-off becomes more pronounced in sparser configurations. Notably, in the most sparse regimes, the post-hoc method can also achieve a lower Type I error rate while maintaining power comparable to the original derandomized approach. In particular, for larger signal amplitudes, our method matches the power of the original procedure while offering improved Type I error control.
Figure 4: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Ren and Barber [21] under a Gaussian linear model. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}$, including the original setting $p_{\text{relevant}}=80$.
Figure 5: Performance comparison between the proposed derandomized post-hoc knockoff procedure and the original derandomized knockoffs method of Ren and Barber [21] under a Logistic linear model. Results are reported across signal amplitudes for varying sparsity levels, given by the number of relevant variables $p_{\text{relevant}}$, including the original setting $p_{\text{relevant}}=60$.
6 A case study using synthetic clinical trial data
To demonstrate the practical utility of our approach and the advantages of post-hoc knockoffs for analysing clinical trial data, we consider a synthetic dataset designed to mimic a real randomized controlled trial. The data were generated using the benchtm R package [30], using Scenario 1 with a continuous outcome and a sample size of $n=1000$. Each patient is described by $p=30$ baseline covariates, a dimensionality that reflects realistic clinical settings where tens of demographic, clinical, laboratory, and lifestyle variables are routinely collected and jointly analysed. The outcome for each patient was generated as a function of covariates $\mathbf{X}=(X_{1},X_{2},\dots,X_{30})$ and treatment assignment $T\in\{0,1\}$ according to
$f(\mathbf{X},T)=s\times\big(0.5\,\mathbb{I}(X_{1}=\text{Y})+X_{11}\big)+T\big(\beta_{0}+\beta_{1}\Phi(20(X_{11}-0.5))\big),$ (34)
where $\mathbb{I}(\cdot)$ denotes the indicator function, $s=2.32$ controls the baseline covariate effect, and $\beta_{0}=0.053$ and $\beta_{1}=0.383$ determine the average treatment effect and the additional effect in the specified subpopulation, respectively. These parameters were tuned to mimic real clinical trial data. A full description of the data-generating mechanism and simulation design is provided in Sun et al. [30].
In this setting, the target variables to be selected by the knockoff procedures are $X_{1}$, $X_{11}$ and the treatment indicator $T$.
We simulate 10 independent RCT datasets following the procedure described above, denoted by R1, …, R10. We examine which variables are selected by the two methods under a nominal FDR level of 0.20. Figure 6 displays the corresponding heatmaps of selected variables.
As shown in the left panel, the original knockoffs method selects five variables only in run R2, and selects none in the remaining nine runs. This implies that, 90% of the time, the method makes no discoveries, leading to frequent failures in detecting true signals. In contrast, the post-hoc method (right panel) selects variables $X_{1}$ and $X_{11}$, both true predictors, in all ten runs, and additionally selects the third relevant variable $T$ in 7 out of 10 runs. The average $\alpha^{ph}$ across the ten runs is 0.32. Moreover, the heatmap shows that the realized FDR is substantially below the nominal target, a trend that is consistently observed across the broader simulations in Section 5.
(a) Original knockoffs heatmap
(b) ph-knockoffs heatmap
Figure 6: Variable selection heatmaps across 10 simulated RCT datasets using the original knockoffs method and the post-hoc knockoffs method. Rows correspond to variables and columns to simulation runs. Blue indicates that a variable was selected in a given run, while white indicates that it was not selected. In this setting, the true relevant variables are $X_{1}$, $X_{11}$ and $T$.
7 Discussion
In this paper, we have shown how the significance level of the knockoff method can be chosen data-dependently while maintaining a reasonable FDR control. Remarkably, this flexibility comes without any price: (1) if the original method does not make any rejections, we can increase the significance level to potentially make rejections and (2) if the original method makes rejections, we can make the same rejections while often reducing (and never enlarging) the significance level, thus increasing the precision. In this way, the results obtained by our post-hoc knockoff procedure are always more informative than the results offered by the original knockoff procedure. This distinguishes our method from existing post-hoc $\alpha$ approaches for other testing problems [8, 14, 35, 6, 5], where the gained flexibility usually reduces the power compared to the standard methods.
Our post-hoc approach solves the well-known issue of the original knockoff method, which requires at least $1/\alpha^{\textnormal{kn}}$ rejections to make any rejection. This makes our framework particularly suitable for settings with small numbers of variables and/or sparse signals.
Furthermore, we integrated the post-hoc idea into derandomized knockoff procedures. While the latter require to choose some arbitrary parameters ($\alpha_{\textnormal{ebh}}$ for FDR and $\eta$ for PFER control) that influences the power and the guarantee of the procedure, we showed that the parameters can be chosen completely data-dependent while maintaining a reasonable control. Ren et al. [22] even mentioned that a data-dependent choice of the parameter $\eta$ would be desirable but is not allowed and thus proposed data-splitting to determine $\eta$ in a data-driven way. Obviously, our approach to choose $\eta$ based on the entire data is much more powerful and practical.
In Section 4.1 we introduced a uniform improvement of the derandomized (ph-)knockoff methods based on the e-Closure Principle. However, due to its high computational requirements, it is difficult to apply in practice. We wonder whether there are computationally efficient short-cuts for this method. This could be explored in future work.